What mass of silver metal is produced from an excess of copper wire added to 4.92g of AgNO3?



1
Expert's answer
2022-05-12T15:27:04-0400

Solution:

The molar mass of AgNO3 is 169.87 g/mol

Therefore,

Moles of AgNO3 = (4.92 g AgNO3) × (1 mol AgNO3 / 169.87 g AgNO3) = 0.02896 mol AgNO3


Balanced chemical equation:

2AgNO3 + Cu → Cu(NO3)2 + 2Ag

According to stoichiometry:

2 mol of AgNO3 produce 2 mol of Ag

Thus, 0.02896 mol of AgNO3 produce:

(0.02896 mol AgNO3) × (2 mol Ag / 2 mol AgNO3) = 0.02896 mol Ag


The molar mass of Ag is 107.868 g/mol

Therefore,

Mass of Ag = (0.02896 mol Ag) × (107.868 g Ag / 1 mol Ag) = 3.1238 g Ag = 3.12 Ag

Mass of Ag = 3.12


Answer: 3.12 grams of silver metal is produced